# Invariant Hilbert spaces of distribution vectors in Lie group representations

### v1

@@ Abstract.
For every unitary irreducible representation of a nilpotent Lie group we prove that the representation Hilbert space is the only nonzero invariant Hilbert space of distribution vectors.
@@
nilpotent Lie group; smooth vector; distribution vector
@@ 1. Introduction
Hilbert subspaces of topological vector spaces in the sense of Laurent Schwartz [Sch64] provide an abstract version of reproducing kernels of function spaces and they thus play an important role in many areas of analysis and representation theory of Lie groups, as shown for instance in [Ne00]. There has been a recent interest in uniqueness of Hilbert subspaces of spaces of tempered distributions that are invariant to modulations and translations, partly motivated by problems related to time-frequency analysis and modulation spaces, see [ToGuMaRa21], [BaMoPi24], [RaToVi25].
In this paper we prove an abstract uniqueness theorem for Lie group representations, which recovers some of the preceding results in the special case of Schrödinger representations of the Heisenberg groups. Specifically, we show that if $\pi\colon G\to{\mathcal{B}}({\mathcal{H}})$ is an irreducible unitary representation of a connected and simply connected nilpotent Lie group $G$, then ${\mathcal{H}}$ is the only continuously embedded Hilbert space into the space of distribution vectors ${\mathcal{H}}_{-\infty}$ which is nonzero and is invariant to the action of $G$ (Theorem 3.1).
Our approach is completely different from the methods from the earlier papers, inasmuch as we rely on some basic results in representation theory of nilpotent Lie groups. Thus, in Corollary 3.4, we provide a new perspective on [RaToVi25, Thm. 0.3].
@@ 2. Preliminaries
Here is our main setting:
  • $G$ is a connected, simply connected nilpotent Lie group
  • $\pi\colon G\to{\mathcal{B}}({\mathcal{H}})$ is an irreducible (continuous) unitary representation;
  • :=${\mathcal{H}}_{\infty}:=\{v\in{\mathcal{H}}\mid\pi(\cdot)v\in\mathcal{C}^{\infty}(G,{\mathcal{H}})\}$ is the Fréchet space of smooth vectors;
  • ${\mathcal{H}}_{-\infty}$ is the space of distribution vectors, i.e., the space of continuous antilinear functionals on ${\mathcal{H}}_{\infty}$, endowed with the $w^{*}$-topology;
  • $\langle\cdot,\cdot\rangle\colon{\mathcal{H}}_{-\infty}\times{\mathcal{H}}_{\infty}\to{\mathbb{C}}$ is the corresponding sesquilinear duality pairing, antilinear in its second variable, which agrees with the scalar product of the Hilbert space ${\mathcal{H}}$;
@@
    $\pi_{\infty}\colon G\to{\mathcal{L}}({\mathcal{H}}_{\infty})\text{ and }\pi_{-\infty}\colon G\to{\mathcal{L}}({\mathcal{H}}_{-\infty})$
    satisfying $\pi_{\infty}(g)=\pi(g)|_{{\mathcal{H}}_{\infty}}$ and $\pi_{-\infty}(g):=(\pi_{\infty}(g^{-1}))^{\prime}$ for every $g\in G$;
  • ${\mathcal{S}}(G)$ is the associative Fréchet algebra of Schwartz functions on $G$ with the convolution product.
@@
${\mathcal{H}}_{\infty}\hookrightarrow{\mathcal{H}}\hookrightarrow{\mathcal{H}}_{-\infty}$ (2.1)
intertwine the group representations $\pi_{\infty}$, $\pi$, and $\pi_{-\infty}$.
Using integration with respect to the Haar measure of $G$, the unitary representation $\pi$ gives rise to a Banach *-algebra representation $\pi\colon L^{1}(G)\to{\mathcal{B}}({\mathcal{H}})$. Moreover, we have $\pi({\mathcal{S}}(G)){\mathcal{H}}_{\infty}\subseteq{\mathcal{H}}_{\infty}$, and this gives rise to the algebra representation
$\pi_{\infty}\colon{\mathcal{S}}(G)\to{\mathcal{L}}({\mathcal{H}}_{\infty}).$
Also, the map
$\pi_{-\infty}\colon{\mathcal{S}}(G)\to{\mathcal{L}}({\mathcal{H}}_{-\infty}),\quad\pi_{-\infty}(\varphi):=(\pi(\varphi^{*}))^{\prime}$
is an algebra representation.
The maps in (2.1) intertwine the representations $\pi_{\infty}$, $\pi|_{{\mathcal{S}}(G)}$, $\pi_{-\infty}$ of the associative algebra ${\mathcal{S}}(G)$ and we have
$\pi_{-\infty}({\mathcal{S}}(G)){\mathcal{H}}_{-\infty}\subseteq{\mathcal{H}}_{\infty}.$ (2.2)
Lemma 2.2.
If ${\mathcal{V}}\subseteq{\mathcal{H}}_{\infty}$ is a ${\mathbb{C}}$-linear subspace satisfying $\pi_{\infty}({\mathcal{S}}(G)){\mathcal{V}}\subseteq{\mathcal{V}}$ then either ${\mathcal{V}}=\{0\}$ or ${\mathcal{V}}={\mathcal{H}}_{\infty}$, that is, the representation $\pi_{\infty}\colon{\mathcal{S}}(G)\to{\mathcal{L}}({\mathcal{H}}_{\infty})$ is algebraically irreducible.
Proof.
Our main result Theorem 3.1 can be regarded as a sharpening of the condition of absence of nontrivial closed invariant subspaces for irreducible unitary representations of nilpotent Lie groups. Before continuing our preparations for its proof, we derive the following simple result that has a similar flavour and may hold an independent interest.
Proposition 2.3.
If ${\mathcal{Y}}\subseteq{\mathcal{H}}_{-\infty}$ is a $w^{*}$-closed linear subspace with $\pi_{-\infty}({\mathcal{S}}(G)){\mathcal{Y}}\subseteq{\mathcal{Y}}$, then we have either ${\mathcal{Y}}=\{0\}$ or ${\mathcal{Y}}={\mathcal{H}}_{-\infty}$.
Proof.
Let ${\mathcal{Y}}^{\perp}:=\{v\in{\mathcal{H}}_{\infty}\mid\langle{\mathcal{Y}},v\rangle=\{0\}\}$, which is a closed linear subspace of ${\mathcal{H}}_{\infty}$. Since ${\mathcal{Y}}\subseteq{\mathcal{H}}_{-\infty}$ is a $w^{*}$-closed linear subspace, we obtain a topological linear isomorphism
${\mathcal{Y}}\simeq({\mathcal{H}}_{\infty}/{\mathcal{Y}}^{\perp})^{\prime}.$ (2.3)
On the other hand, the hypothesis $\pi_{-\infty}({\mathcal{S}}(G)){\mathcal{Y}}\subseteq{\mathcal{Y}}$ implies $\pi_{\infty}({\mathcal{S}}(G)){\mathcal{Y}}^{\perp}\subseteq{\mathcal{Y}}^{\perp}$ hence, by Lemma 2.2, we have either ${\mathcal{Y}}^{\perp}=\{0\}$ or ${\mathcal{Y}}^{\perp}={\mathcal{H}}_{\infty}$. Now the assertion follows by (2.3). ∎
The next lemma is a more precise form of [Ho77, Cor. 3.4.2]. We give below its short proof.
Lemma 2.4.
Let $\pi$ be as above and $\sigma\colon G\to{\mathcal{B}}({\mathcal{X}})$ be another unitary representation of $G$. Assume that the linear map $T\colon{\mathcal{H}}_{\infty}\to{\mathcal{X}}$ satisfies
Proof.
Let $v_{1},v_{2}\in{\mathcal{H}}_{\infty}\setminus\{0\}$ be such that $\langle v_{1},v_{2}\rangle=0$. The orthogonal projection $P:=\langle\cdot,v_{1}\rangle{v_{1}}$ belongs to ${\mathcal{B}}({\mathcal{H}})_{\infty}$ by [BeBe10, Ex. 3.1]. By [Ho77, Thm. 3.4], there is $\varphi\in{\mathcal{S}}(G)$ with $P=\pi(\varphi)$, hence $P=P^{*}=P^{2}=\pi(\psi)$ where $\psi:=\varphi^{*}\ast\varphi$. Then, by hypothesis, $\sigma(\psi)T=TP=\langle\cdot,v_{1}\rangle Tv_{1}$, hence
$\sigma(\psi)Tv_{1}=Tv_{1}\;\;\text{and}\;\;\sigma(\psi)Tv_{2}=0.$
It follows that $Tv_{1}$, respectively $Tv_{2}$, are eigenvectors for the eigenvalues $1$, respectively 00, of the self-adjoint operator $\sigma(\psi)$, and thus they are orthogonal.
Therefore we have obtained that for $v_{1},v_{2}\in{\mathcal{H}}_{\infty}$,
Let now $v,w\in{\mathcal{H}}_{\infty}$ be arbitrary with $\|v\|=\|w\|=1$. Set $v_{2}=w-\langle w,v\rangle v$. Then $\langle v,v_{2}\rangle=0$ hence, by (2.4), $\langle Tv,Tv_{2}\rangle=0$ , that is,
$\langle Tv,Tw\rangle=\langle v,w\rangle\|Tv\|^{2}.$ (2.5)
Interchanging the roles of $w$ and $v$ we have obtained that $\langle v,w\rangle\|Tv\|^{2}=\langle v,w\rangle\|Tw\|^{2}$. It follows that
$\|Tv\|=\|Tw\|\quad\text{if }\langle v,w\rangle\neq 0.$
Letting $\epsilon\to 0$, we get that
$\|Tv\|=\|Tw\|\quad\text{for every}\quad v,w\in{\mathcal{H}}_{\infty},\;\|v\|=\|w\|=1.$
Thus $\lambda_{0}:=\|Tv\|$ is independent on $v\in{\mathcal{H}}_{\infty}$ with $\|v\|=1$. Then by (2.5) we get
$\langle Tv,Tw\rangle=\lambda_{0}^{2}\langle v,w\rangle$ (2.6)
[…]
@@ 3. The main result
Let $\pi\colon G\to{\mathcal{B}}({\mathcal{H}})$ be an irreducible unitary representation of a connected, simply connected nilpotent Lie group $G$. Let ${\mathcal{X}}\subseteq{\mathcal{H}}_{-\infty}$ be a ${\mathbb{C}}$-linear subspace endowed with the structure of a complex Hilbert space satisfying the following hypotheses:
@@
    the inclusion map ${\mathcal{X}}\hookrightarrow{\mathcal{H}}_{-\infty}$ is continuous;
@@
    $\pi_{-\infty}(G){\mathcal{X}}\subseteq{\mathcal{X}}$;
@@
    for every $g\in G$ the operator $\pi_{\mathcal{X}}(g):=\pi_{-\infty}(g)|_{\mathcal{X}}\colon{\mathcal{X}}\to{\mathcal{X}}$ is an isometry.
Then either ${\mathcal{X}}=\{0\}$ or ${\mathcal{X}}={\mathcal{H}}$.
For the proof of the above theorem we still need the following lemma.
@@
We assume ${\mathcal{X}}\neq\{0\}$. For every $v\in{\mathcal{H}}_{\infty}$ we have $\overline{\langle\cdot,v\rangle}\in({\mathcal{H}}_{-\infty})^{\prime}$, hence $\overline{\langle\cdot,v\rangle}|_{\mathcal{X}}\in{\mathcal{X}}^{\prime}$ by (H1). Since ${\mathcal{X}}$ is a complex Hilbert space, there exists a uniquely determined vector $\iota(v)\in{\mathcal{X}}$ satisfying
@@
$\iota\colon{\mathcal{H}}_{\infty}\to{\mathcal{X}}.$
The equality (3.1) shows that $\iota$ is the adjoint of the inclusion map ${\mathcal{X}}\hookrightarrow{\mathcal{H}}_{-\infty}$. The range of the operator $\iota$ is dense in ${\mathcal{X}}$ since the above inclusion map is injective. In fact, if $x_{0}\in{\mathcal{X}}$ satisfies $x_{0}\perp\iota({\mathcal{H}}_{\infty})$ in ${\mathcal{X}}$ then, by (3.1), we have $\langle x_{0},v\rangle=0$ for every $v\in{\mathcal{H}}_{\infty}$, that is, $x_{0}=0$.
In order to prove that $\iota$ is a continuous operator, one can use the Closed Graph Theorem since ${\mathcal{H}}_{\infty}$ is a Fréchet space and ${\mathcal{X}}$ is a Hilbert space. Specifically, if $\lim\limits_{n\to\infty}v_{n}=v$ in ${\mathcal{H}}_{\infty}$ and $\lim\limits_{n\to\infty}\iota(v_{n})=x$ in ${\mathcal{X}}$, then $\iota(v)=x$ by (3.1).
For every $g\in G$, $v\in{\mathcal{H}}_{\infty}$, and $x\in{\mathcal{X}}$ we have
$\langle x,\iota(\pi_{\infty}(g)v)\rangle_{\mathcal{X}}$ $=\langle x,\pi_{\infty}(g)v\rangle=\langle\pi_{-\infty}(g^{-1})x,v\rangle=\langle\pi_{\mathcal{X}}(g^{-1})x,v\rangle$
$=\langle\pi_{\mathcal{X}}(g^{-1})x,\iota(v)\rangle_{\mathcal{X}}=\langle x,\pi_{\mathcal{X}}(g)\iota(v)\rangle_{\mathcal{X}}$
@@
$(\forall g\in G,v\in{\mathcal{H}}_{\infty})\quad\iota(\pi_{\infty}(g)v)=\pi_{\mathcal{X}}(g)\iota(v).$ (3.2)
@@
Since $\iota\colon{\mathcal{H}}_{\infty}\to{\mathcal{X}}$ is a continuous linear operator, and the mapping $G\ni g\mapsto\pi_{\infty}(g)v\in{\mathcal{H}}_{\infty}$ is smooth as $\pi\colon G\to{\mathcal{B}}({\mathcal{H}})$ is a continuous representation (cf. [Po72, Prop. 1.2]), it also follows by (3.2) that for every $x\in\iota({\mathcal{H}}_{\infty})\subseteq{\mathcal{X}}$ the mapping $\rho_{\mathcal{X}}^{x}\colon G\to{\mathcal{X}}$, $\pi_{\mathcal{X}}^{x}(g):=\pi_{\mathcal{X}}(g)x$, is smooth, and in particular continuous. We proved above that $\iota({\mathcal{H}}_{\infty})$ is dense in ${\mathcal{X}}$, hence by, (H3), Lemma 3.2 is applicable and ensures that $\pi_{\mathcal{X}}\colon G\to{\mathcal{B}}({\mathcal{X}})$ is a continuous unitary representation, and moreover $\iota({\mathcal{H}}_{\infty})\subseteq{\mathcal{X}}_{\infty}$. Then, using (3.2) once again, we obtain
$(\forall\varphi\in{\mathcal{S}}(G))\quad\iota\circ\pi_{\infty}(\varphi)=(\pi_{\mathcal{X}})_{\infty}(\varphi)\circ\iota\colon{\mathcal{H}}_{\infty}\to{\mathcal{X}}_{\infty}.$
This implies by Lemma 2.4 that there exist $\lambda_{0}\in{\mathbb{R}}$ and an isometry $\iota_{0}\colon{\mathcal{H}}_{\infty}\to{\mathcal{X}}$ with $\iota=\lambda_{0}\iota_{0}$. Since $\iota$ is has dense range in ${\mathcal{X}}$ and thus is non-trivial when ${\mathcal{X}}\neq\{0\}$, $\lambda_{0}\neq 0$ and we may assume that $\lambda_{0}>0$, by replacing $\iota_{0}$ by $-\iota_{0}$ if necessary.
Using again that the range of $\iota$ is dense in ${\mathcal{X}}$ it follows that $\iota_{0}\colon{\mathcal{H}}_{\infty}\to{\mathcal{X}}$ extends by continuity to a unitary operator $\iota_{0}\colon{\mathcal{H}}\to{\mathcal{X}}$ that intertwines the representations $\pi$ and $\pi_{\mathcal{X}}$. In particular, the unitary representation $\pi_{\mathcal{X}}$ is irreducible.
@@
On the other hand, we recall from (2.2) that $\pi_{-\infty}({\mathcal{S}}(G)){\mathcal{H}}_{-\infty}\subseteq{\mathcal{H}}_{\infty}$, in particular $\pi_{-\infty}({\mathcal{S}}(G)){\mathcal{X}}\subseteq{\mathcal{H}}_{\infty}$, while the hypothesis (H2) along with the continuity of the representation $\pi_{\mathcal{X}}(\cdot)=\pi_{-\infty}(\cdot)|_{\mathcal{X}}$ implies by (2.2) again that $\pi_{-\infty}({\mathcal{S}}(G)){\mathcal{X}}\subseteq{\mathcal{X}}_{\infty}$. Thus
$\pi_{-\infty}({\mathcal{S}}(G)){\mathcal{X}}\subseteq{\mathcal{X}}_{\infty}\cap{\mathcal{H}}_{\infty}$
and then the property ${\mathcal{X}}\neq\{0\}$ shows that ${\mathcal{X}}_{\infty}\cap{\mathcal{H}}_{\infty}\neq\{0\}$. As above, by the hypothesis (H2) again we also have $\pi_{\infty}({\mathcal{S}}(G))({\mathcal{X}}_{\infty}\cap{\mathcal{H}}_{\infty})\subseteq{\mathcal{X}}_{\infty}\cap{\mathcal{H}}_{\infty}$. Then, by Lemma 2.2 applied for both irreducible representations $\pi$ and $\pi_{\mathcal{X}}$, we obtain ${\mathcal{H}}_{\infty}={\mathcal{X}}_{\infty}\cap{\mathcal{H}}_{\infty}={\mathcal{X}}_{\infty}$, that is, ${\mathcal{H}}_{\infty}={\mathcal{X}}_{\infty}$.
By another application of Lemma 2.4, the identity map ${\rm id}_{{\mathcal{H}}_{\infty}}\colon{\mathcal{H}}_{\infty}\to{\mathcal{X}}$ extends to ${\mathcal{H}}$ and its extension is a scalar multiple of an isometry. Since ${\mathcal{H}}_{\infty}$ is dense in ${\mathcal{H}}$, and ${\mathcal{X}}_{\infty}$ is dense in ${\mathcal{X}}$, we get that ${\mathcal{H}}\subset{\mathcal{X}}$, this extension is a multiple of an unitary operator and it is nothing else than ${\rm id}_{{\mathcal{H}}}\colon{\mathcal{H}}\to{\mathcal{X}}$. That is, ${\mathcal{X}}={\mathcal{H}}$ as sets, and the scalar products of these Hilbert spaces are a suitable scalar multiple of each other. ∎
@@
In the setting of Theorem 3.1, if we assume (H2), then (H3) is equivalent to the following, apparently weaker, condition:
@@
In fact, clearly (H3$\Rightarrow$(H4) with $\|\cdot\|^{\prime}_{\mathcal{X}}=\|\cdot\|^{\prime}_{\mathcal{X}}$ and $C=1$. For the converse implication, let us assume (H4) and denote by $G_{d}$ the group $G$ endowed with its discrete topology. Since $G_{d}$ is a nilpotent group, it is solvable hence amenable. See e.g., [Gr69, page 9]. Then, by [Gr69, Thm. 3.4.1] and its proof, one can define an equivalent Hilbert space norm $\|\cdot\|_{\mathcal{X}}$ on ${\mathcal{X}}$ for which the representation $\pi_{\mathcal{X}}\colon G_{d}\to{\mathcal{B}}({\mathcal{X}})$ is unitary, that is, (H3) holds true.

### v2

@@ Abstract.
For every unitary irreducible representation of a Lie group we prove that the representation Hilbert space is the only nonzero invariant Hilbert space of distribution vectors.
@@
nilpotent Lie group; smooth vector; reproducing kernel
@@ 1. Introduction
Hilbert subspaces of topological vector spaces in the sense of Laurent Schwartz [Sch64] provide an abstract version of reproducing kernels of function spaces and they thus play an important role in many areas of mathematical physics, analysis, and representation theory of Lie groups, as shown for instance in [Ne00]. There has been a recent interest in uniqueness of Hilbert subspaces of spaces of tempered distributions that are invariant to modulations and translations, partly motivated by problems related to time-frequency analysis and modulation spaces, see [ToGuMaRa21], [BaMoPi24], [RaToVi25].
In this paper we prove an abstract uniqueness theorem for Lie group representations, which recovers some of the preceding results in the special case of Schrödinger representations of the Heisenberg groups. Specifically, we show that if $\pi\colon G\to{\mathcal{B}}({\mathcal{H}})$ is an irreducible unitary representation of a Lie group $G$, then $\mathcal{H}$} is the only continuously embedded Hilbert space into the space of distribution vectors $\mathcal{H}$$^{-\infty}$ which is nonzero and is invariant to the action of $G$ (Theorem 3.1). In section 4 we give a version of that result replacing the action of the group by the action of the Schwartz algebra on a nilpotent Lie group.
As explained in [ToGuMaRa21], the minimality properties of the Feichtinger algebra provided the initial motivation for uniqueness results on Hilbert spaces that are invariant to translations and modulations. In that spirit, one used the realization of the Hilbert space as $L^{2}({\mathbb{R}}^{d})$, which facilitated the techniques of distribution theory and Hermite functions. Our approach is completely different, inasmuch as we rather rely on some basic results on representations of Lie groups and thus, in Corollary 3.4, we provide a new perspective on [RaToVi25, Thm. 0.3].
@@ 2. Preliminaries
We need a version of Schur’s lemma for unbounded intertwining operators.
Definition 2.1.
Let $G$ be a locally compact group and $\pi\colon G\to{\mathcal{B}}({\mathcal{H}}_{j})$, $j=1,2$, two unitary representations of the group $G$. An unbounded operator $S$ from ${\mathcal{H}}_{1}$ to ${\mathcal{H}}_{2}$ with domain $D(S)\subseteq{\mathcal{H}}_{1}$ is an intertwining operator for the representations $\pi_{1}$ and $\pi_{2}$ if for every $g\in G$ and $x\in D(S)$ we have $\pi(g)x\in D(S)$ and
$S\pi_{1}(g)x=\pi_{2}(g)Sx.$
Lemma 2.2.
Let $G$ be a locally compact group and $\pi\colon G\to{\mathcal{B}}({\mathcal{H}}_{j})$, $j=1,2$, two unitary representations of the group $G$ with $\pi_{1}$ irreducible. Assume that $S$ is non-zero closable intertwining operator for $\pi_{1}$ and $\pi_{2}$. Then the closure $\overline{S}$ of $S$ is a bounded operator from ${\mathcal{H}}_{1}$ to ${\mathcal{H}}_{2}$ and there is $\lambda>0$ such that $\lambda\overline{S}\colon{\mathcal{H}}_{1}\to{\mathcal{H}}_{2}$ is an isometry.
Proof.
This is a direct consequence of [Ne00, Lemma II.2.6 and Thm. II.2.8]. ∎
In the present paper, unless otherwise mentioned, we use the following notation.
  • $G$ is a Lie group;
  • $\pi\colon G\to{\mathcal{B}}({\mathcal{H}})$ is a (continuous) unitary representation;
  • ${\mathcal{H}}^{\infty}:=\{v\in{\mathcal{H}}\mid\pi(\cdot)v\in\mathcal{C}^{\infty}(G,{\mathcal{H}})\}$ is the Fréchet space of smooth vectors;
  • $\mathcal{H}$$^{-\infty}$ is the space of distribution vectors, i.e., the space of continuous antilinear functionals on $\mathcal{H}$$^{\infty}$, endowed with the $w^{*}$-topology;
  • $\langle\cdot,\cdot\rangle\colon{\mathcal{H}}^{-\infty}\times$$\mathcal{H}$$^{\infty}\to{\mathbb{C}}$ is the corresponding sesquilinear duality pairing, antilinear in its second variable, which agrees with the scalar product of the Hilbert space $\mathcal{H}$};
@@
[…]
    satisfying $\pi^{\infty}(g)=\pi(g)|_{{\mathcal{H}}^{\infty}}$ and $\pi^{-\infty}(g):=(\pi^{\infty}(g^{-1}))^{\prime}$ for every $g\in G$;
  • $C_{0}^{\infty}(G)$ is the space of compactly supported smooth functions on $G$.
@@
$\mathcal{H}$$^{\infty}\hookrightarrow{\mathcal{H}}\hookrightarrow$$\mathcal{H}$$^{-\infty}$ (2.1)
intertwine the group representations $\pi^{\infty}$, $\pi$, and $\pi^{-\infty}$.
Using integration with respect to a Haar measure of $G$, the unitary representation $\pi$ gives rise to a Banach *-algebra representation $\pi\colon L^{1}(G)\to{\mathcal{B}}({\mathcal{H}})$. Moreover, we have $\pi(C_{0}^{\infty}(G)){\mathcal{H}}^{\infty}\subseteq{\mathcal{H}}^{\infty}$ with dual map
$\pi^{-\infty}\colon C_{0}^{\infty}(G)\to{\mathcal{L}}({\mathcal{H}}^{-\infty}),\quad\pi^{-\infty}(\varphi):=(\pi(\varphi^{*}))^{\prime}.$
Then
$\pi^{-\infty}(C_{0}^{\infty}(G)){\mathcal{H}}^{-\infty}={\mathcal{H}}^{\infty}.$ (2.2)
by [DM78, Thm. 3.1 and 3.2].
@@ 3. The main result
Let $\pi\colon G\to{\mathcal{B}}({\mathcal{H}})$ be an irreducible unitary representation of a Lie group $G$. Let $\mathcal{X}$$\subseteq{\mathcal{H}}^{-\infty}$ be a $\mathbb{C}$}-linear subspace endowed with the structure of a complex Hilbert space satisfying the following hypotheses:
@@
    the inclusion map $\mathcal{X}$$\hookrightarrow{\mathcal{H}}^{-\infty}$ is continuous;
@@
    $\pi^{-\infty}(G){\mathcal{X}}\subseteq{\mathcal{X}}$;
@@
    for every $g\in G$ the operator $\pi_{\mathcal{X}}(g):=\pi^{-\infty}(g)|_{\mathcal{X}}\colon{\mathcal{X}}\to{\mathcal{X}}$ is an isometry.
Then either ${\mathcal{X}}=\{0\}$ or $\mathcal{X}$$\mathcal{H}$}.
For the proof of the above theorem we need the following lemma.
@@
We assume ${\mathcal{X}}\neq\{0\}$. For every $v\in{\mathcal{H}}^{\infty}$ we have $\overline{\langle\cdot,v\rangle}\in({\mathcal{H}}^{-\infty})^{\prime}$, hence $\overline{\langle\cdot,v\rangle}|_{\mathcal{X}}\in{\mathcal{X}}^{\prime}$ by (H1). Since $\mathcal{X}$} is a complex Hilbert space, there exists a uniquely determined vector $\iota(v)\in{\mathcal{X}}$ satisfying
@@
$\iota\colon{\mathcal{H}}^{\infty}\to{\mathcal{X}},$
such that
$\iota(v)=v\quad\text{for }v\in{\mathcal{H}}^{\infty}\cap{\mathcal{X}}.$
The equality (3.1) shows that $\iota$ is the adjoint of the inclusion map $\mathcal{X}$$\hookrightarrow{\mathcal{H}}^{-\infty}$. The range of the operator $\iota$ is dense in $\mathcal{X}$} since the above inclusion map is injective. In fact, if $x_{0}\in{\mathcal{X}}$ satisfies $x_{0}\perp\iota({\mathcal{H}}^{\infty})$ in $\mathcal{X}$} then, by (3.1), we have $\langle x_{0},v\rangle=0$ for every $v\in{\mathcal{H}}^{\infty}$, that is, $x_{0}=0$.
In order to prove that $\iota$ is a continuous operator, one can use the closed graph theorem since $\mathcal{H}$$^{\infty}$ is a Fréchet space and $\mathcal{X}$} is a Hilbert space. Specifically, if $\lim\limits_{n\to\infty}v_{n}=v$ in $\mathcal{H}$$^{\infty}$ and $\lim\limits_{n\to\infty}\iota(v_{n})=x$ in $\mathcal{X}$}, then $\iota(v)=x$ by (3.1).
For every $g\in G$, $v\in{\mathcal{H}}^{\infty}$ and $x\in{\mathcal{X}}$ we have
$\langle x,\iota(\pi^{\infty}(g)v)\rangle_{\mathcal{X}}$ $=\langle x,\pi^{\infty}(g)v\rangle=\langle\pi^{-\infty}(g^{-1})x,v\rangle=\langle\pi_{\mathcal{X}}(g^{-1})x,v\rangle$
$=\langle\pi_{\mathcal{X}}(g^{-1})x,\iota(v)\rangle_{\mathcal{X}}=\langle x,\pi_{\mathcal{X}}(g)\iota(v)\rangle_{\mathcal{X}},$
@@
$(\forall g\in G,v\in{\mathcal{H}}^{\infty})\quad\iota(\pi^{\infty}(g)v)=\pi_{\mathcal{X}}(g)\iota(v).$ (3.2)
@@
Since $\iota\colon{\mathcal{H}}^{\infty}\to{\mathcal{X}}$ is a continuous linear operator and the mapping $G\ni g\mapsto\pi^{\infty}(g)v\in{\mathcal{H}}^{\infty}$ is smooth, as $\pi\colon G\to{\mathcal{B}}({\mathcal{H}})$ is a continuous representation (cf. [Po72, Prop. 1.2]), it also follows by (3.2) that, for every $x\in\iota({\mathcal{H}}^{\infty})\subseteq{\mathcal{X}}$, the mapping $\rho_{\mathcal{X}}^{x}\colon G\to{\mathcal{X}}$, $\pi_{\mathcal{X}}^{x}(g):=\pi_{\mathcal{X}}(g)x$, is smooth and in particular continuous. We proved above that $\iota({\mathcal{H}}^{\infty})$ is dense in $\mathcal{X}$}, hence by (H3), Lemma 3.2 is applicable and ensures that $\pi_{\mathcal{X}}\colon G\to{\mathcal{B}}({\mathcal{X}})$ is a continuous unitary representation. Moreover, (3.2) shows that $\iota({\mathcal{H}}^{\infty})\subseteq{\mathcal{X}}^{\infty}$, hence
${\mathcal{H}}^{\infty}\cap{\mathcal{X}}\subseteq{\mathcal{X}}^{\infty}.$ (3.3)
@@
On the other hand, we recall from (2.2) that $\pi^{-\infty}(C_{0}^{\infty}(G))){\mathcal{H}}^{-\infty}={\mathcal{H}}^{\infty}$; in particular, $\pi_{{\mathcal{X}}}(C_{0}^{\infty}(G)){\mathcal{X}}\subseteq{\mathcal{H}}^{\infty}$. Then Dixmier-Malliavin’s theorem [DM78, Thm. 3.2] for the continuous representation $\pi_{\mathcal{X}}$ shows that $\mathcal{X}$$^{\infty}\subseteq{\mathcal{H}}^{\infty}$ which, along with (3.3) gives
${\mathcal{X}}^{\infty}={\mathcal{H}}^{\infty}\cap{\mathcal{X}}.$ (3.4)
The subspace $\mathcal{X}$$^{\infty}$ is then contained in $\mathcal{H}$} and dense in $\mathcal{X}$} . Moreover since it is invariant under $\pi(g)$, $g\in G$, and the representation $\pi\colon G\to{\mathcal{B}}({\mathcal{H}})$ is irreducible, it follows that $\mathcal{X}$$^{\infty}$ is dense in $\mathcal{H}$}, as well.
We now regard the operator $\iota$ above as an unbounded operator defined on $\mathcal{H}$} with domain $\mathcal{H}$$^{\infty}$. A vector $x\in{\mathcal{X}}$ belongs to the domain of the adjoint $\iota^{*}$ of $\iota$ if and only if the map $v\mapsto\langle x,\iota(v)\rangle_{\mathcal{X}}=\langle x,v\rangle$ is continuous in the norm of $\mathcal{H}$}. Since the duality pairing $\langle\cdot,\cdot\rangle$ agrees with the scalar product of $\mathcal{H}$}, we obtain that $\mathcal{X}$$^{\infty}$ is contained in the domain of $\iota^{*}$. It follows that $\iota^{*}$ is densely defined, therefore $\iota$ is a closable operator. Using now (3.2) and Lemma 2.2 we get that there are an isometry $\iota_{0}\colon{\mathcal{H}}\to{\mathcal{X}}$ and a constant $\lambda_{0}>0$ such that $\iota=\lambda_{0}\iota_{0}|_{{\mathcal{H}}^{\infty}}$. As the range of $\iota$, hence of $\iota_{0}$, is dense in $\mathcal{X}$}, it follows that the operator $\iota_{0}\colon{\mathcal{H}}\to{\mathcal{X}}$ is unitary. Moreover, since $\iota|_{{\mathcal{X}}^{\infty}}$ is the identity map, and $\mathcal{X}$$^{\infty}$ is dense both in $\mathcal{X}$} and $\mathcal{H}$}, we get that ${\mathcal{H}}={\mathcal{X}}$ as sets and the scalar products of these Hilbert spaces are a suitable scalar multiple of each other. ∎
@@
In the setting of Theorem 3.1, we denote by $G_{d}$ the group $G$ endowed with its discrete topology. If we assume that $G_{d}$ is amenable and (H2) holds true, then (H3) is equivalent to the following, seemingly weaker, condition:
@@
[…]
We note that if the Lie group $G$ is connected, the condition that $G_{d}$ is amenable is equivalent with $G$ being solvable. (See [Pa88, Thm. 3.9].)